OpenStudy (anonymous):

The height-to-arm-length ratio of a tyrannosaurus is 20 feet to 3 feet. A human is 8 1/2 heads tall. Use this ratio to determine the arm length of a human if the human were built like a tyrannosaurus.

OpenStudy (dumbcow):

is that 8.5 feet tall

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you solve it?

OpenStudy (anonymous):

can someone sove this?

OpenStudy (dumbcow):

ok set up the ratio like this two fractions set equal to each other T_height/T_armLength = H_height/H_armLength 20/3 = 8.5/x cross multiply 20x = 3*8.5 x=25.5/20

OpenStudy (anonymous):

is that star multiply?

OpenStudy (dumbcow):

yes